OpenStudy (miguel980):

The cost of a ticket to enter a local amusement park is $50. The park charges an additional $2.50 per ride. A function that describes this is C(r)=50+2.5r, where C(r) is the total cost and r is the number of rides purchased. For the inverse function, the independent variable is the total cost and the dependent variable is the number of rides purchased. Which of these describes the inverse function's domain?

OpenStudy (miguel980):

A. \[C \ge 50\] B.\[C>0\] C. Domain: \[r \ge 0\] where r is a whole number D. Domain: \[C \ge 50 \] in whole number multiples of 2.50

OpenStudy (anonymous):

I think it is D but im not sure